Assume that the risk-free rate is 6% and the market risk premium is 3%.

What is the required return for the overall stock market? Round your answer to two decimal places.
  %

What is the required rate of return on a stock with a beta of 1.4? Round your answer to two decimal places.
  %
Answer #1

market risk premium=market rate-risk free rate

market rate=6+3=9%

Required rate=risk free rate+beta*(market rate-risk free rate)

=6+1.4*(9-6)

=10.2%